# Simultaneous Localization and 3D-Semi Dense Mapping for Micro Drones Using Monocular Camera and Inertial Sensors

### v1

@@ I Introduction
Classical visual SLAM systems, such as ORB-SLAM [mur2015orb, 1], rely on sparse feature matching, geometric triangulation, and global optimization to estimate camera pose and map structure. While effective for pose estimation, these methods often produce sparse maps ”3D point” that lack structural richness, making them less suitable for tasks requiring dense 3D understanding or interaction, While Voxel map  [1] offers improved retrieval and reasoning over visibility than ORB-SLAM  [mur2015orb], it remains fundamentally sparse storing only a few landmark points per voxel and still lacks the structural and geometric richness required for dense perception and navigation. Moreover, their reliance on computationally intensive operations like loop closure  [mur2015orb, 2] (a process that detects when the camera revisits a scene to improve accuracy) and global bundle adjustment [3] (an optimization technique that refines camera poses and 3D points simultaneously) can make real-time deployment on embedded systems impractical. Furthermore, Structure from Motion (SfM) is a foundational technique that reconstructs 3D structures from multiple images, serving as a precursor to SLAM. SfM begins by detecting and extracting features using methods such as SIFT [4], SURF [5], ORB [6], or learned features [7]. These features are matched across images, with outliers removed through geometric verification techniques like RANSAC [8]. Camera positions are estimated from these correspondences, and 3D points are triangulated to build a sparse reconstruction. Bundle adjustment then refines the entire structure by minimizing reprojection errors [3]. Despite its effectiveness, SfM faces similar challenges to SLAM in real-time applications, especially on resource-limited platforms, while Edge SLAM [9] has been proposed as an alternative that aims to overcome these limitations by generating semi-dense maps but remains computationally heavy due to reliance on global optimization techniques such as loop closure. Additionally, their tracking loss function based on optical flow introduces inherent noise, which can degrade accuracy and hinder real-time performance on limited hardware.

### v2

@@ I Introduction
Classical visual SLAM systems, such as ORB-SLAM [1, 2], rely on sparse feature matching, geometric triangulation, and global optimization to estimate camera pose and map structure. While effective for pose estimation, these methods often produce sparse maps ”3D point” that lack structural richness, making them less suitable for tasks requiring dense 3D understanding or interaction, While Voxel map  [2] offers improved retrieval and reasoning over visibility than ORB-SLAM  [1], it remains fundamentally sparse storing only a few landmark points per voxel and still lacks the structural and geometric richness required for dense perception and navigation. Moreover, their reliance on computationally intensive operations like loop closure  [1, 3] (a process that detects when the camera revisits a scene to improve accuracy) and global bundle adjustment [4] (an optimization technique that refines camera poses and 3D points simultaneously) can make real-time deployment on embedded systems impractical. Furthermore, Structure from Motion (SfM) is a foundational technique that reconstructs 3D structures from multiple images, serving as a precursor to SLAM. SfM begins by detecting and extracting features using methods such as SIFT [5], SURF [6], ORB [7], or learned features [8]. These features are matched across images, with outliers removed through geometric verification techniques like RANSAC [9]. Camera positions are estimated from these correspondences, and 3D points are triangulated to build a sparse reconstruction. Bundle adjustment then refines the entire structure by minimizing reprojection errors [4]. Despite its effectiveness, SfM faces similar challenges to SLAM in real-time applications, especially on resource-limited platforms, while Edge SLAM [10] has been proposed as an alternative that aims to overcome these limitations by generating semi-dense maps but remains computationally heavy due to reliance on global optimization techniques such as loop closure. Additionally, their tracking loss function based on optical flow introduces inherent noise, which can degrade accuracy and hinder real-time performance on limited hardware.